# General Hamiltonian description of nonreciprocal interactions

### v3

@@ IV Equivalence between Hamiltonian Monte-Carlo and Langevin Dynamics
We provide evidence for the equivalence of the Langevin and Glauber descriptions in numerical simulations. In Fig. 3, we compare (a) the average magnetization $\langle m\rangle{=}\sqrt{(\sum\cos\theta_{i})^{2}{+}\left(\sum\sin\theta_{i}\right)^{2}}/L^{2}$ and (b) the specific heat, $C_{v}{=}\left(\left\langle H_{SS}^{2}\right\rangle{-}\left\langle H_{SS}\right\rangle^{2}\right)/\left(LT\right)^{2}$, obtained as the variance of the energy in the original spin layer [56], as a function of temperature, for constrained Glauber dynamics (blue) vs. Langevin dynamics (green). We initialize the system in an ordered state, $\theta_{i}{=}0,\varphi_{i}{=}\pi$, with initial energy of the original system $H_{SS}{=}3L^{2}$. Consistently with previous works [39, 29, 40, 41], our data indicates the existence of a phase transition between an ordered phase at low temperature and disordered phase at high temperature [Fig. 3(a)]. This critical behavior is corroborated by the specific heat $C_{v}$, which exhibits a peak at the phase transition point [see Fig. 3(b)]. Importantly, both the Langevin and the constrained Monte-Carlo simulations exhibit the same critical temperature, $T_{c}/J=0.78\pm 0.02$ (Langevin) and $T_{c}/J=0.79\pm 0.03$ (constrained Glauber), for the phase transition, within statistical uncertainty, see SI [43]. By contrast, a Monte-Carlo simulation using the Hamiltonian (3) without including the auxiliary spins $\mathbf{a}_{i}$ produces a critical temperature of $T_{c}/J{\approx}1.72$ [43, 57] (not shown).
@@
Unlike Gauber dynamics based on the selfish energy [30], our constrained-Hamiltonian Monte-Carlo update rule results in the same configuration-space density for the steady state as the Langevin equation. Indeed, the critical temperatures for the phase transition estimated from the Langevin and the selfish-energy approach do not agree in general [compare green and red curves in Fig. 3], although for some parameters the values may agree within statistical uncertainty. By ensuring the equivalence of the Langevin and the constrained Monte-Carlo dynamics, our Hamiltonian embedding resolves this long-standing issue in the study of vision-cone interactions [45].
@@
We emphasize that, whereas we used a specific model to demonstrate the equivalence of the Monte-Carlo dynamics with the original nonreciprocal Langevin equation, the concept is generally applicable to other nonreciprocal systems, as we show in [43].
@@ V Hamiltonian engineering for nonreciprocal systems
The Hamiltonian framework for nonreciprocal interactions that we propose has a natural symplectic structure which, besides applications in Monte-Carlo simulations, enables the use of canonical transformations to study the physics of nonreciprocal systems. As a concrete example, here we demonstrate the application of Floquet engineering to modify the properties of nonreciprocal systems using high-frequency periodic driving [42, 58, 59]. This enables one to reach regimes of interest that might be inaccessible in experiments due to the limited tunability of parameters.
Consider again the vision-cone spins on a square lattice and assume that one has no control over the magnitude and sign of the interaction strength; such situations are abundant in experimental systems. We will now show how high-frequency periodic driving allows us to effectively modify the magnitude and sign of the interaction strength. Let us subject the system to periodic driving, motivated by a condensed-matter setting  [^2] The XY model arises naturally in the description of Josephson junction arrays. When subject to an external AC magnetic field, the latter can be modeled by a time-dependent vector potential which couples to the phase, $\theta_{i}$, of the superconducting order parameter [77]., following the equations of motion
@@
We model the external periodic drive along the $x$-direction with frequency $\omega$ and amplitude $A_{ij}{=}aA_{0}\cos\psi_{ij}$, where the angle $\psi_{ij}$ is defined in Fig. 2a, and $a$ is the lattice constant [61, 62, 63]. We will show that the periodic drive allows us to effectively suppress the spin interactions along the $x$-direction only [Fig. 4a] while keeping the vision-cone interaction along the $y$-direction — an instance of Floquet engineering known as dynamical localization [42, 58, 59]. In systems with limited control over the interaction strength, this strategy would allow experiments to study the fate of the phase transition, as the dimensionality of the system is effectively changed from $d=2$ (a square lattice) to $d=1$ (a collection of uncoupled chains).
The Hamiltonian formalism proves convenient to analyze the driven system due to its symplectic structure. Adding the auxiliary spin $\mathbf{a}_{j}$, we construct the time-periodic Hamiltonian embedding
@@
with $\left\{\theta_{j},\varphi_{i}\right\}{=}\delta_{ij}$. To derive an effective description at high driving frequencies, we apply the Floquet-Magnus expansion $H_{F}{=}H_{F}^{(0)}{+}\mathcal{O}(\omega^{-1})$ [42]; the leading-order Floquet Hamiltonian is the time-average
[…]
@@
where $\mathcal{J}_{0}\left(A_{ij}\right)$ is the Bessel function of the first kind.
@@
Figure 4: The Hamiltonian embedding enables the use of Floquet techniques to analyze driven nonreciprocal interacting systems. (a) Schematic illustration of the periodically driven XY spins (left panel, Eq. (V)) and the effective time-averaged interactions (right panel, Eq. (11)). Thin black dashed lines indicate suppression of interactions in the $x$-direction. (b) Average magnetization $\langle m\rangle$ obtained from Langevin dynamics as a function of temperature $T$ for different drive amplitude strengths $A_{0}$. As $A_{0}$ increases, the transition temperature decreases; as predicted by the time-averaged Floquet Hamiltonian, when $A_{0}{=}z_{n}a^{-1}$ (see text), the system no longer exhibits a long-range ordered phase. We use $10^{5}$ iterations with a timestep of $\delta t=0.01$; $\omega{=}20\pi J$, and $\Psi=7\pi/8$. (c-d) Average magnetization $\langle m\rangle$ obtained from Langevin dynamics as a function of time $t$ for two drive frequencies: $\omega{=}\pi J$ (c) and $\omega{=}20\pi J$ (d), for $T{=}0.5J$. As the frequency increases, the dynamics corresponding to the equations of motion derived from $H_{F}^{(0)}$ (orange) become closer to the exact periodically driven Langevin dynamics (purple). Both show a significant deviation from the Langevin dynamics of the non-driven system from Eq. (1) (green). Parameters: $A_{0}{=}2a^{-1}$, and $L{=}64$.
@@
To verify the validity of the Floquet Hamiltonian description, we study the average magnetization as a function of time $t$ for two driving frequencies in Fig. 4(c,d). The deviation between the results from the exact Langevin equations of motion (purple) and the effective equations of motion derived from $H_{F}^{(0)}$ (red) decreases as the frequency increases [from Fig. 4(c) to Fig. 4(d)]; moreover, the drive modifies the behavior significantly (compare red and green curves).
@@
By tuning the amplitude $A_{0}$ of the drive close to the zero of the Bessel function, $\mathcal{J}_{0}(z_{n}){=}0$, one can suppress spin interactions in the $x$-direction, thereby realizing a dimensional crossover from a two-dimensional lattice to a set of one-dimensional chains [Fig. 4a]. Figure 4(b) shows the average magnetization obtained from simulations of the spin dynamics (9) in the high-frequency regime as a function of temperature $T$. As we gradually increase $A_{0}$, the transition temperature $T_{c}$ decreases. To leading order in $\omega^{-1}$, when $A_{0}{=}z_{n}a^{-1}$, the system decouples into a set of independent 1D XY-chains and no longer exhibits an ordered phase at any finite reservoir temperature. The behavior in Fig. 4(b) is consistent with our analysis using the Floquet Hamiltonian $H_{F}^{(0)}$, since the dimensional crossover occurs precisely at the zeroes of the Bessel function. Finally, note that, even though periodic drives typically lead to enhanced energy absorption, the heating rates are either power-law or exponentially suppressed with increasing the drive frequency $\omega$ [64].
@@ VI Discussion and Outlook
According to Liouville’s theorem, without constraint, the Hamiltonian embedding gives rise to nondissipative dynamics in the embedding phase space, even if the original equations of motion are dissipative. This feature is brought about by the symplectic structure inherent to the Hamiltonian embedding (e.g., in the vision-cone XY spins that we discussed, $\theta$ and $\varphi$ are conjugate variables). By contrast, on the submanifold defined by the constraint, the dynamics are dissipative. This apparent contradiction is resolved by noticing that the constrained submanifold is a set of measure zero in the embedding phase space which does not contribute to phase-space volume. However, the mechanism by which dissipation emerges from constrained Hamiltonian dynamics presents an intriguing direction for future research, closely related to a formulation of Hamiltonian dynamics for nonconservative systems [56].
When evaluated on the constraint (5), the total Hamiltonian $H$ vanishes identically; therefore, $H$ does not describe the energy of the original nonreciprocal system, but merely provides a generator for its dynamics. Moreover, the ground-state manifold of $H$ need not obey the constraint. These unusual features arise as a result of the original nonreciprocal system being intrinsically out of equilibrium.
@@
Crucially, our construction works for pairwise interactions, and may not apply to arbitrary non-potential systems [cf. Fig. 1]; a generalization to velocity-dependent (e.g., Lorentz-like) forces is an interesting direction for future work. Unlike in systems where non-reciprocity arises as a result of projecting/tracing out bath degrees of freedom (e.g., a chemical concentration field) that mediate the interactions [1], we emphasize that here the auxiliary degree of freedom does not play the role of a bath for the original one; instead, the projecting/tracing out procedure is replaced by the constraint. The framework is generic and applies to both inertial and dissipative dynamics featuring nonreciprocal pairwise interactions.
@@
We discussed two applications using XY spins on a square lattice with vision-cone interactions: (i) We proved the equivalence between Glauber dynamics based on the Hamiltonian embedding and the original nonreciprocal Langevin dynamics. In particular, we showed that both procedures yield the same critical temperature for the transition between the ferromagnetic and non-magnetized phases. (ii) We make use of the inherent symplectic structure of the embedding to derive an effective Hamiltonian for periodically driven vision-cone XY spins using Floquet theory; this allows us to interpret the disappearance of ferromagnetic order when varying the drive amplitude as a dimensional crossover. These results suggest exciting possibilities for experiments to apply Floquet engineering techniques [42, 58, 59] to investigate regimes in parameter space that are otherwise inaccessible in the nonreciprocal system.
@@
We note that the Hamiltonian embedding does not introduce new physics content per se beyond what is already captured by the equations of motion defining the original nonreciprocal system. However, it enables the study of nonreciprocal systems using analytical and numerical tools from equilibrium or non-equilibrium physics that require a Hamiltonian structure. For instance, one can now use Monte Carlo algorithms that shortcut long transients of Langevin simulations and lead directly to the steady state. Likewise, one can apply various canonical transformations (e.g., Schrieffer-Wolff-like transformations such as the Birkhoff-Gustavson normal form [34, 35, 36]) to derive nonreciprocal models with effective degrees of freedom. An illustrative example is the Floquet-Magnus expansion for periodically driven systems: higher-order terms in the inverse drive frequency contain effective spin-spin interactions [42]. Note that whereas it is possible to derive effective stroboscopic equations of motion in the high-frequency regime directly from the equations of motion to any order in the inverse frequency $\omega^{-1}$ by using two-times perturbation theory [66], the Hamiltonian formulation has the advantage of providing a systematic model-agnostic algorithmic procedure [64]. This formalism can prove useful to analyze the behavior of time-crystalline nonreciprocal states [67].

### v4

@@ IV.1 Nonequilibrium steady states
In Fig. 3, we compare (a) the average magnetization $\langle m\rangle{=}\left\langle\sqrt{(\sum_{i}\cos\theta_{i})^{2}{+}\left(\sum_{i}\sin\theta_{i}\right)^{2}}/L^{2}\right\rangle$ and (b) the specific heat, $C_{v}{=}\left(\left\langle H_{SS}^{2}\right\rangle{-}\left\langle H_{SS}\right\rangle^{2}\right)/\left(LT\right)^{2}$, obtained as the variance of the energy in the original spin layer [32], as a function of temperature, for constrained Glauber dynamics (blue) vs. Langevin dynamics (green). We initialize the system in an ordered state, $\theta_{i}{=}0$, with initial energy of the original system $H_{SS}{=}3L^{2}$. Consistent with previous works [20, 51, 62, 25], our data indicates the existence of a phase transition between an ordered phase at low temperature and disordered phase at high temperature [Fig. 3(a)]. This critical behavior is corroborated by the specific heat $C_{v}$, which exhibits a peak at the phase transition point [see Fig. 3(b)]. Importantly, both the Langevin and the constrained Monte-Carlo simulations exhibit the same critical temperature, $T_{c}/J=0.79\pm 0.04$, for the phase transition, within statistical uncertainty, see SI [1]. By contrast, a Monte-Carlo simulation using the Hamiltonian (3) without the effect of the auxiliary spins $\mathbf{a}_{i}$ produces a critical temperature of $T_{c}/J{\approx}1.72$ [1, 8] (not shown).
@@
Unlike Gauber dynamics based on the selfish energy [5], our constrained-Hamiltonian Monte-Carlo update rule results in the same configuration-space density for the nonequilibrium steady state as the Langevin equation. Indeed, the critical temperatures for the phase transition estimated from the Langevin and the selfish-energy approach do not coincide in general [compare green and red curves in Fig. 3], although for some parameters the values may agree within statistical uncertainty. By ensuring the equivalence of the Langevin and the constrained Monte-Carlo dynamics, our Hamiltonian embedding resolves this long-standing issue in the study of vision-cone interactions [67].
IV.2 Nonstationary states
Figure 4: Nonstationary state of XY spins with chase-and-run interactions. (a) XY spins on a square lattice with nonreciprocal interactions along the $x$-direction, and reciprocal interactions along the $y$-direction. The x=1x{=}1 column used to define the order parameter $O(t)$ is marked in red. (b) Stripe pattern used as an initial steady state, which moves perpetually to the left due to the nonreciprocal interactions (see Video). (c)/(d) Long-time evolution of the real/imaginary part of the stripe order parameter $\langle O(t)\rangle$ averaged over an ensemble of $\mathcal{N}{=}100$ independent realizations of the noisy dynamics. The data show excellent agreement between the Langevin [dashed line, top $x$-axis] and the constrained Glauber simulations [solid line, bottom $x$-axis] with transition rates obtained from the Hamiltonian embedding. The time axis is rescaled by the linear system size $L$. The parameters are $_{\rightarrow}/J{=}1, J/J$$_{\leftarrow}/J{=}{-}0.99$, $J_{\text{rec}}/J{=}10$, T/J=0.08T/J{=}0.08, $\Delta{=}0.05$, and $L=200$, with the scaling factor $\alpha(\Delta,T){=}2.37{\times}10^{-3}$ extracted numerically [1].
We now show that the Hamiltonian embedding also describes nonreciprocal systems that exhibit long-time nonstationary states, such as sustained oscillations, corresponding to the breaking of time-translation symmetry.
Consider the dynamics of XY spins on a square lattice following Eq. (1), but with nonreciprocal interactions $J_{\leftarrow}{\neq}J_{\rightarrow}$ along the $x$-direction (independent of the spin orientation), and reciprocal interactions $J_{\text{rec}}$ along the $y$-direction, see Fig. 4(a). The initial state is the vertical stripe pattern shown in Fig. 4(b). Whenever $J_{\leftarrow}{\approx}{-}J_{\rightarrow}$, the spins exhibit a chase-and-run dynamics [30] and the pattern moves to the left (for $J_{\rightarrow}{>}0$) at a constant speed (see Video). In the thermodynamic limit, this nonstationary state propagates perpetually due to an interplay between noise and many-body interactions [1], breaking continuous time-translation symmetry. These sustained dynamics do not have an equilibrium analogue, and hence present an ideal testbed for the validity of our constrained Glauber dynamics.
Following the procedure in Sec. III, we calculate a Hamiltonian embedding for this model (Methods); then, using Eq. (6), we find the expression for the energy difference $\Delta E(\theta,\theta^{\prime})$, cf. Eq. (27).
Figure 4(c-d) shows the dynamics of the complex-valued order parameter $O(t)=L^{-1}\sum_{y=1}^{L}e^{i\theta_{1,y}(t)},$ which characterizes the average orientation of the spins within the x=1x{=}1 column of the lattice [shown in red in Fig. 4(a)]. Then, with periodic boundary conditions, the real and imaginary parts of $\langle O(t)\rangle$ exhibit temporal oscillations as a result of the collective spin rotation at the interface between the stripes, which moves due to the nonreciprocal interactions. As anticipated from our analytical theory, we find an excellent agreement between the Langevin and Monte-Carlo dynamics at all times, including transients [1]. This provides numerical evidence that Glauber dynamics with transition rates determined from the Hamiltonian embedding capture correctly the nonstationary-state behavior of nonreciprocal systems.
@@ IV.2 Nonstationary states
We emphasize that, while above we used two specific models to demonstrate the equivalence between Monte-Carlo dynamics and the original nonreciprocal Langevin equation, the Hamiltonian embedding applies to generic nonreciprocal systems, as we show in [1].
@@ V Hamiltonian engineering for nonreciprocal systems
Whereas the Glauber dynamics in the previous section rely on the embedding Hamiltonian to define transition rates, they do not make use of the underlying symplectic (or phase-space) structure, as Monte-Carlo simulations only require an energy function. The symplectic structure of the embedding Hamiltonian enables the use of canonical transformations to analyze the physics of nonreciprocal systems, e.g., in identifying sets of strongly and weakly interacting degrees of freedom which define fast and slow variables, or in determining conserved quantities. Its existence is also inherently related to the existence of a Hamiltonian as a generator of dynamics.
As a concrete example of leveraging the symplectic structure provided by the embedding Hamiltonian, here we demonstrate the application of Hamiltonian engineering to modify the properties of nonreciprocal systems using high-frequency periodic driving [18, 36, 29]. To this end, we consider a hypothetical experiment on a system of XY spins interacting nonreciprocally on a square lattice (Fig. 5a). Hamiltonian engineering allows one to tune the interaction strength independently along the different lattice directions, which may enable one to explore regimes not accessible otherwise. As an example, below we show how to use high-frequency periodic driving to modify the interactions and achieve a dimensional crossover, whereby a square lattice of spins with vision-cone interactions is effectively turned into a one-dimensional chain (Fig. 5a).
To define the drive, note that the XY model arises naturally in the description of Josephson junction arrays; when subject to an external AC magnetic field, the field can be modeled by a time-dependent vector potential $A_{ij}$ which couples to the phase difference $\theta_{i}{-}\theta_{j}$ of the superconducting order parameter [57]. While the discussion here is not bound to a specific experimental system, this setting illustrates concretely how a periodic drive can modify the interaction term when it is not possible to directly tune its coupling strength $J_{ij}$. In the presence of the drive, the equations of motion read as
@@
We model the external periodic drive along the $x$-direction with frequency $\omega$ and amplitude $A_{ij}{=}aA_{0}\cos\psi_{ij}$, where the angle $\psi_{ij}$ is defined in Fig. 2a, and $a$ is the lattice constant [26, 50, 3].
We now demonstrate that the periodic drive enables us to effectively suppress the spin interactions along the $x$-direction only [Fig. 5a] while keeping the vision-cone interaction along the $y$-direction   [^3] In the literature on quantum systems [18, 36, 29], this phenomenon is known as dynamical localization, since an external drive is used to suppress the transfer of spin-state population between neighboring lattice sites, controlled by the corresponding spin-spin interaction strength. In the case of nonreciprocal interactions, the periodic drive leads to weakening and eventually complete decoupling along the $x$-bonds of the lattice. – an instance of Floquet engineering [18, 36, 29]. In systems with limited control over the interaction strength, this strategy would allow experiments to study the fate of the phase transition, as the dimensionality of the system is effectively changed from $d=2$ (a square lattice) to $d=1$ (a collection of uncoupled chains). However, the Floquet engineering framework uses a Hamiltonian, which is not directly available in nonreciprocal systems.
The Hamiltonian formalism from Sec. III proves convenient to analyze the driven system due to its symplectic structure. Adding the auxiliary spin $\mathbf{a}_{j}$, we construct the time-periodic Hamiltonian embedding
@@
with $\left\{\theta_{j},\varphi_{i}\right\}{=}\delta_{ij}$. To derive an effective description at high driving frequencies, $J_{ij}/\omega{\ll}1$, we apply the Floquet-Magnus expansion $H_{F}{=}H_{F}^{(0)}{+}\mathcal{O}(\omega^{-1})$ [18]; the leading-order Floquet Hamiltonian is the time-average
$H_{F}^{(0)}$ ${=}$ ${-}\sum_{\langle ij\rangle}\mathcal{J}_{0}\left(A_{ij}\right)J_{ij}(\theta_{i})\left[\cos\left(\theta_{i}{-}\theta_{j}\right){+}\cos\left(\varphi_{i}{-}\theta_{j}\right)\right]$ (10)
@@
where $\mathcal{J}_{0}\left(A_{ij}\right)$ is the Bessel function of the first kind. Higher-order terms in the Floquet-Magnus expansion can be readily obtained in a straightforward way using this formalism; they control the validity of the approximation.
@@
Figure 5: The Hamiltonian embedding enables the use of Floquet techniques to analyze driven nonreciprocal interacting systems. (a) Schematic illustration of the periodically driven XY spins (left panel, Eq. (V)) and the effective time-averaged interactions (right panel, Eq. (10)). Thin black dashed lines indicate suppression of interactions in the $x$-direction. (b) Average magnetization $\langle m\rangle$ obtained from Langevin dynamics as a function of temperature $T$ for different drive amplitude strengths $A_{0}$. As $A_{0}$ increases, the transition temperature decreases; as predicted by the time-averaged Floquet Hamiltonian, when $A_{0}{=}z_{\ast}a^{-1}$ (see text), the system no longer exhibits a long-range ordered phase. We use $10^{5}$ iterations with a timestep of $\delta t=0.01$; $\omega{=}20\pi J$, and $\Psi=7\pi/8$. (c-d) Average magnetization $\langle m\rangle$ obtained from Langevin dynamics as a function of time $t$ for two drive frequencies: $\omega{=}\pi J$ (c) and $\omega{=}20\pi J$ (d), for T=0.5JT{=}0.5J. As the frequency increases, the dynamics corresponding to the equations of motion derived from $H_{F}^{(0)}$ (orange) become closer to the exact periodically driven Langevin dynamics (purple). Both show a significant deviation from the Langevin dynamics of the non-driven system from Eq. (1) (green). Parameters: $A_{0}{=}2a^{-1}$, and L=64L{=}64.
@@
To verify the validity of the Floquet Hamiltonian description in the nonreciprocal setting, we study the average magnetization as a function of time $t$ for two drive frequencies in Fig. 5(c,d). The deviation between the results from the exact Langevin equations of motion (purple) and the effective equations of motion derived from $H_{F}^{(0)}$ (red) decreases as the frequency increases [from Fig. 5(c) to Fig. 5(d)]; moreover, the drive modifies the behavior significantly (compare red and green curves).
@@
By tuning the amplitude $A_{0}$ of the drive close to a zero of the Bessel function, $\mathcal{J}_{0}(z_{\ast}){=}0$, one can suppress spin interactions in the $x$-direction, thereby realizing a dimensional crossover from a two-dimensional lattice to a set of one-dimensional chains [Fig. 5a]. Figure 5(b) shows the average magnetization obtained from simulations of the spin dynamics (8) in the high-frequency regime as a function of temperature $T$. As we gradually increase the drive amplitude $A_{0}$, the transition temperature $T_{c}$ decreases. To leading order in $\omega^{-1}$, when $A_{0}{=}z_{\ast}a^{-1}$, the system decouples into a set of independent 1D XY-chains and no longer exhibits an ordered phase at any finite reservoir temperature. The behavior in Fig. 5(b) is consistent with our analysis using the Floquet Hamiltonian $H_{F}^{(0)}$, since the dimensional crossover occurs precisely at the zeroes of the Bessel function. Finally, note that, even though periodic drives typically lead to enhanced energy absorption, the heating rates are either power-law or exponentially suppressed with increasing the drive frequency $\omega$ [55].
In summary, Floquet-Hamiltonian engineering based on the framework from Sec. III provides enhanced control over model parameters in nonreciprocal systems, enabling access to wider regions of their phase diagrams and allowing one to probe dimensional crossovers. More broadly, the Hamiltonian embedding – of which computing the Floquet Hamiltonian is one application – unlocks generic Hamiltonian methods for nonreciprocal systems. We anticipate its symplectic structure to similarly enable the transfer of other techniques to nonreciprocal systems.
@@ VI Discussion and Outlook
Crucially, our construction works for pairwise interactions, and may not apply to arbitrary non-potential systems [cf. Fig. 1]; a generalization to velocity-dependent (e.g., Lorentz-like) forces is presented in [1]. Unlike in systems where nonreciprocity arises as a result of projecting/tracing out bath degrees of freedom that mediate the interactions (e.g., a chemical concentration field) [16], we emphasize that here the auxiliary degree of freedom does not play the role of a bath for the original one; instead, the projecting/tracing out procedure is replaced by the constraint. The framework is generic and applies to both inertial and dissipative dynamics featuring nonreciprocal pairwise interactions.
Our Hamiltonian embedding advances a century-old effort of using auxiliary degrees of freedom to describe generic nonconservative forces. First attempts date back to Bateman [10], where a Lagrangian was constructed for a dissipative oscillator; however, these early results cannot be generalized beyond harmonic oscillators as they neglect the dynamical constraint from Eq. (5). Markovian embeddings have also been frequently used to describe and simulate non-Markovian systems with memory kernels [73, 45, 24]. More recently, Galley presented a general formalism to describe nonconservative forces by formulating Hamilton’s principle as an initial value problem [32]: it requires doubling the degrees of freedom and relies on an educated guess for the interaction term that couples them. This approach produces the correct nonreciprocal equations of motion, when supplemented with our Hamiltonian (2). A key contribution of our work is the generic construction of the interaction energy between the original and the auxiliary degrees of freedom for any pairwise nonreciprocal interaction; previous literature has not discussed nonreciprocal interactions within a Hamiltonian framework. Critically, our approach differs from Galley’s in the implementation: Galley imposes an equality condition on the degrees of freedom at the final time, from which the so-called physical limit [32] arises by setting the boundary terms in the variational principle to vanish; the conceptual difference in our approach is the constraint that we impose at the initial time, which is then automatically preserved by the dynamics at all times.
@@
We discussed two applications using XY spins on a square lattice with vision-cone interactions: (i) We proved the equivalence between Glauber dynamics based on the Hamiltonian embedding and the original nonreciprocal Langevin dynamics. In particular, we showed that both procedures yield the same critical temperature for the transition between the ferromagnetic and non-magnetized phases. These results open the door to investigating the nature of this and similar phase transitions in large systems using fast Monte-Carlo algorithms. (ii) We make use of the inherent symplectic structure of the embedding to derive an effective Hamiltonian for periodically driven vision-cone XY spins using Floquet theory; this allows us to interpret the disappearance of ferromagnetic order when varying the drive amplitude as a dimensional crossover. These results suggest exciting possibilities for experiments to apply Floquet engineering techniques [18, 36, 29] to investigate regimes in parameter space that are otherwise inaccessible in the nonreciprocal system.
@@
We note that the Hamiltonian embedding does not introduce new physics content per se beyond what is already captured by the equations of motion defining the original nonreciprocal system. However, it enables the study of nonreciprocal systems using analytical and numerical tools from equilibrium or nonequilibrium physics that require a Hamiltonian structure. For instance, one can now use Monte-Carlo algorithms that shortcut long transients of Langevin simulations and lead directly to the steady state. Numerical techniques based on the embedding may require keeping track of twice the number of original degrees of freedom; however, the numerical complexity of any algorithm remains unchanged by adding the auxiliary degrees of freedom (which effectively doubles the system size). Likewise, one can apply various canonical transformations (e.g., Schrieffer-Wolff-like transformations such as the Birkhoff-Gustavson normal form [14, 40, 70]) to identify fast and slow effective degrees of freedom in nonreciprocal models. An illustrative example is the Floquet-Magnus expansion for periodically driven systems: higher-order terms in the inverse drive frequency contain effective spin-spin interactions [18]. Note that whereas it is possible to derive effective stroboscopic equations of motion in the high-frequency regime directly from the equations of motion to any order in the inverse frequency $\omega^{-1}$ by using two-times perturbation theory [79], the Hamiltonian formulation has the advantage of providing a systematic model-agnostic algorithmic procedure. This formalism can prove useful to analyze the behavior of time-crystalline nonreciprocal states [41].